# General Riemann Integration on Partially Ordered Net Convergence Structures

### v1

@@ Abstract
In this article, we propose a general theory of integration of the Riemann type with respect to arbitrary measures and functions, connected by a continuous bilinear product, with values in abstract Riesz spaces endowed with a general convergence given by nets. This covers both the topological and order based convergences in the literature. We then show that this integral satisfies most of the common order based properties of the objects that comprises integration theory. By establishing a generalized notion of summability on Riesz spaces and an integral build upon countable partitions of the base space and the summability theory, we then shown some uniform convergence theorems for the refereed integrals. An application of these theorems is made to prove various equivalences concerning the Lebesgue, for which we give a brief survey, and Riemann type integrals in partially ordered spaces defined previously. We finish the article with a possible way of classifying general integration procedures defined in abstract convergence structures, and pose some open problems based on them.
@@
1 Introduction and Historical Context
Since the pioneering work of Riemann [1] on trigonometric functions and integration, a myriad of different types of integrals from the said "Riemann type" have been proposed. Putting aside the much studied Henstock-Kurzweil case, we also have, in the case of partially ordered spaces and with much prominence, the Riemann-Stieltjes integral of [2], [3] and [4], the Birkhoff type integral of [5], the SL, Abstract Derivation Basis and Choquet integrals of [6], [7], [8] and [9], as well as many others that can be gauged from these references and the surveys of [10], [11] and [12]. Therefore, generalized Riemann integration of Henstock-Kurzweil and Choquet types, and the vector-scalar partially ordered case of Birkhoff are well developed from both the general theory, and applications, point of views.
On the other hand, little has been done in generalizing Kolmogorov [13] all encompassing project of integration theory based on Riemann sums to partially ordered spaces. In fact, the author has only been able to localize the works of [14] and [15], which deals with a type of $S^{*}$ integration in convergence groups with examples of the partially ordered case, and the papers of Ionescu-Tulcea [16], [17], [18], which treats the above mentioned Kolmogorov $S^{*}$-integral in the partially ordered Riesz space case, with various examples and simultaneous consideration of the topological case. We also cite the Saks type integrals of [19] and [20] that fit in the general Kolmogorovian framework and will be commented on.
Considering this context and the lack of unification of the integration procedures cited above, we propose, in the present work, to define a general integral based on nets of Riemann sums - using the works of [21], [22], [23], [24] in the topological case - that can be defined in general algebraic structures, equipped with a partial order, and endowed with a specific type of general convergence structure in the sense of [25], called a net convergence structure and recently studied by [26] and [27], but already present in some aspects in [28]. Various additivity and convergence properties of this integration procedure will show that it preserves most of the common properties found in the aforementioned works.
Comparison of this general definition of integral with various types of order convergence based Riemann integrals in the literature (see again [12] for a comprehensive review and bibliography) is also a focus of the present work, and we will show that most of them, as well as Saks type integration procedures, fall withing our framework. Finally, we discuss the equivalence/containment of this general Riemann integral with sequential/net-type approximation Lebesgue type integrals of [29], [30], [31], [32] and others, and propose a general classification and brief bibliographic review of schemes of integration structures available in the literature of partially ordered spaces-valued integration. For information about the topological, algebraic and measure-theoretic concepts that will be used in the context of integration theory in this article, we cite, besides de already given references, [33], [34], [35], [36], [37], [38], [39], [40] and, specially for integration, the more recent publications of [41], [42] and [43].
2 Preliminaries on Notation, Partially Ordered Spaces and Convergence
@@
In in this work, we will, in general, denote partially ordered spaces of any type with uppercase Latin letters $X,Y,Z$, etc. As a rule, all partial orders in the refereed spaces will be denoted by $\leq$, following the custom of classical and recent works on spaces equipped with partial orders, as [35], [36], [44], and [45], as we will not work with more than one partial order, in a fixed space, at a time (therefore, minimizing the possibility of confusion). Differentiation between order notations will be made only when necessary. We will also suppose that every partially ordered space below is Archimedean - for more details, see [36].
As we will be dealing with different types of integrals in this work, we identify them by an superscript consisting of an upper case letter with a direct summary of the name, e.g. $\int^{R}$ for the Riemann integral. The only exception will be the main integral of Section 3, which will be denoted without superscript.
@@
The last point on notation: for a general set $A$, following [46], we denote by $\mathcal{F}(A)$ the collection of all its finite subsets (which is a directed set by inclusion $\subseteq$). We also have the following definition (see, for example, [47]):
@@
We say that a pair $(\Omega,\mathcal{H})$ consisting of an arbitrary set and a collection of its subsets is a Paved Space if $\emptyset\in\mathcal{H}$.
For the measure and integration results below, we shall need a variety of different convergence notions used in the literature, sometimes not equivalently defined in it, and even not equivalent in general. In this case, comparison of our results with other works should be done considering in detail the definitions below. We usually adopt the most general one to avoid problems of compatibility, and point out equivalences (and differences) wherever possible.
We begin by the following definition given in [30] and [45], which originally is due to Vulikh [35], called $(o_{3})$-convergence:
@@
Let $X$ be a partially ordered set and $I$, $C$, and $D$ directed sets. We say that a net $\{x_{i}\}_{i\in I}$ in $X$ $(o_{3})$-converges to $x$ in $X$, which we denote by:
@@
if there is an increasing net $\{z_{c}\}_{c\in C}$ and a decreasing net $\{y_{d}\}_{d\in D}$ in $X$ such that:
@@
Sometimes we may get a simplification by obtaining increasing and decreasing nets in the definition above with same indexing directed set, and then get the following definition of [27] and [26] (see also [45]):
@@
Let $X$ be a partially ordered set and $I$, $C$ directed sets. We say that a net $\{x_{i}\}_{i\in I}$ in $X$ order converges, or $(o)$-converges, to $x$ in $X$, which we denote by:
@@
if there is an increasing net $\{z_{c}\}_{c\in C}$ and a decreasing net $\{y_{c}\}_{d\in C}$ in $X$ such that:
@@
Definition 2.4 (Bounded Net).
Let $X$ be a partially ordered space and $\{x_{i}\}_{i\in I}$ a net in $X$. We say that the net is bounded if and only if there exists an $y\in X$:
@@
The following characterization of order convergent nets will be useful later (see [48] and [49]):
@@
Let $X$ be a lattice ordered group. Then, a net $\{x_{i}\}_{i\in I}$ is $(o)$-convergent to $x\in X$ if and only if it is bounded and:
@@
For lattice ordered groups, we have the following result of [30] and [50], which is also given in [27] and [26] for Riesz spaces:
@@
Let $X$ be a lattice ordered group and $(I,\alpha)$ a directed set. A net $\{x_{i}\}_{i\in I}$ is order convergent in $X$ if and only if it is $o_{3}$-convergent, which occurs if and only if there is a decreasing net $\{y_{d}\}_{d\in D}$ in $X$ such that $y_{d}\downarrow 0$ and for every $d\in D$ there exists $i^{*}\in I$ such that:
@@
Most of the time we will work with objects less general than lattice ordered groups, e.g. Riesz spaces, in such a way that all of the previous convergence definitions, including the one given in the proposition above, are equivalent. Also, we can show, at least for sequences, that the following result is true (see [45]):
@@
Let $X$ be a lattice ordered group. A sequence $\{x_{n}\}_{n\in\mathbb{N}}$ in $X$ is $(o_{3}),o$-convergent to $x$ in $X$ if and only if there is a non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ in $X$ such that $u_{n}\downarrow 0$ and:
@@
The sequential nature of the result above is essential, as "tail" type convergence results for nets are not usually not true. For more about this subject, see page 184 of [45].
For various proofs in the analytical theory of partially ordered spaces, one hard to work with property of the previous convergences is that $\varepsilon$-type arguments have no analogue. This complicates some proofs related to Cauchy-type conditions and countable sequence and subsequence arguments. To remediate this fact, a different type of convergence, usually spoken as a type of "regulator" or $(D)$-convergence was created, and plays a major role in classical texts, such as [37] and [36], as well as in the vast literature in integration and convergence theory in partially ordered spaces, as given by the surveys of [12], [11], and the recent works of [27] and [26], and their collaborators.
In this case we introduce the following structure associated with $(D)$-convergence, as in [11]:
@@
Let $X$ be a partially ordered space. We say that a double sequence $\{a_{ij}\}_{(i,j)\in\mathbb{N}\times\mathbb{N}}$ [^1] From now on, except when strictly needed, we will represent a regulator only as $\{a_{ij}\}_{ij}$. in $X$ is a $(D)$-sequence, or regulator, if $a_{ij}\geq a_{ij+1}$ for every $i,n\in\mathbb{N}$ and:
@@
Let $X$ be a lattice ordered group and $I$ a directed set. We say that a net $\{x_{i}\}_{i\in I}$ in $X$ is $(D)$-convergent to an element $x\in X$ if there exists a $(D)$-sequence $\{a_{ij}\}$ in $X$ such that for every $\varphi\in\mathbb{N}^{\mathbb{N}}$ there exists $i_{0}\in I$ such that:
@@
In general, we have the following relations between this convergence and the previously defined ones (see Chapter 2 of [11]):
@@
Let $X$ be an lattice ordered group. Then, a $(o)$-convergent sequence $\{x_{n}\}_{n\in\mathbb{N}}$ in $X$ is $(D)$-convergent.
In general, the reverse statement of convergence is not equivalent - see, for example, [36]. The result for nets is also not, in general, true - see again Chapter 2 of [11].
As indicated earlier, $(D)$-convergence can give, in some cases, a substitute for $\varepsilon\setminus 2^{n}$-type arguments in terms of (one version of) the Fremlin Lemma (see [10]):
@@
A detailed proof of this important result can be found in [51].
We also have the next concept that will play a pivotal role in proving some convergence theorems for the integrals defined in this work. Originally, at least in integration theory, it is due to Matthes [52] and Maitland-Wright [53].
Definition 2.7 (Weak $\sigma$-Distributivity).
Let $X$ be a lattice ordered group. We say that $X$ is weakly $\sigma$-distributive if and only if:
for every $(D)$-sequence $\{a_{ij}\}$ in $X$.
As an interesting consequence (which we shall use later in the context of integration), we have (see [11]):
Proposition 2.5.
Let $X$ be an Archimedean and weakly $\sigma$-distributive lattice ordered group. Then, a $(D)$-convergent sequence $\{x_{n}\}_{n\in\mathbb{N}}$ in $X$ is $(o)$-convergent.
We finish this introductory section with the next essential definition, which can be found in most books about partially ordered spaces (see, for example, the classic of [35]):
Definition 2.8 (Dedekind Completeness).
For further properties relating to Dedekind completeness, see Chapter 4 of [36].
3 Partially Ordered Net Convergence Structures and Riemann Integration
From the point of view of integration theory, specially considering the breadth and generality of Riemann integration and the various type of convergences in partially ordered spaces defined in the last section, its interesting to consider convergence structures, compatible with the algebraic operations involved, that generalize and unify them - the partially ordered and topological cases being only specific examples of this general structure.
This idea is not new, and actually resurfaced independently in various places in the integration theory literature, such as [14], [12], [40], [54], [15] [55] and [56], all which cites the partially ordered case as the main motivation, or as a particular interesting example. We will leave the development of the general theory of such integration procedures for general spaces, not necessarily partially ordered, for another article, and focus primarily on the case of net convergence structures, in the sense of [27] and [26], with further axioms stated by [40] (see also [12]) and [39], which are tailor made for our work on Riemann integral, in conjunction with the construction of [22] and [24].

### v2

@@
Abstract Integration in Net Convergence Structures
Alexandre Reggiolli Teixeira riccimath123@gmail.com, a163407@dac.unicamp.br
@@ Abstract
In this article, we propose a general theory of integration of the Riemann and Lebesgue types with respect to arbitrary measures and functions, connected by a continuous bilinear product, with values in abstract vector spaces endowed with a convergence structure given by nets. This covers both the topological and order based convergences in the literature. We then show that this integral satisfies most of the common properties of the objects that comprises integration theory. By establishing a generalized notion of summability on vector spaces with a convergence structure and an integral built upon countable partitions of the base space, we then stablish some uniform, monotone and dominated convergence theorems for the refereed integrals, as well as a non-topological or order based Henstock Lemma and a general convergence theorem based on the notion of conjugated lattice seminorms. An application of these theorems is made to prove various equivalences concerning the Lebesgue, for which we give a brief survey, Saks and Riemann type integrals in partially ordered and topological vector spaces presented in the literature, for which we also make a thorough review. We finish the article with a possible way of classifying general integration procedures defined in abstract convergence structures, and pose some open problems based on them.
@@
1 Introduction, Historical Context and Some Notation.
Since the pioneering works on integration theory of Pettis (1938), Bochner (1933), Birkhoff (1935), Bartle (1956) and others, in the beginning to mid of the last century, a myriad of different integration procedures, with varying generality with respect to algebraic, set and convergence theoretic notions, have been proposed, as can be gauge from the vast bibliography in the classical work of Henstock (1991). In almost all cases, these works are constructed on the basis of algebraic and convergence operations on a fixed vector (group/semigroup) space, with the approximation procedure necessary for each integration operation being taken on a fixed topology or order structure. Then, elementary properties of these integrals, such as additivity, preservation of order (isotonocity), subset integration, and others, are obtained using the special approximation procedures - Riemann sums or Lebesgue-type limit arguments - in each case. Further, and more complicated, continuity properties of these are usually constructed by restricting the space of integrable functions or set functions involved, with a special use of convergence sequences of functions or other objects. These are the basic elements of integration theory outlined by Dinculeanu (2000) in his vast treatise on bilinear vector integration, but as can be seen from the recent surveys of Boccuto and Dimitriou (2015) and Boccuto et al. (2012b), the partially ordered spaces construction is no different.
In contrast to these early (and modern) trends, we propose to work with general archetypes of integration procedures, such as Lebesgue and Riemann ones, to find general patterns in the definition/approximation procedures of each of these types (of integration) in such a way that the properties cited in the previous paragraphs are natural consequences, in a very general context, of simple and unifying principles having to do with the process itself, and not deriving from simplifying assumptions related to topology or measure theory.
The first step in this program is to notice that many of the most elementary of these properties, such as additivity of integration or isotonicity in the presence of partial orders, or even the definition of the integration procedure itself, are not dependent on any type of topological or order based convergence, being common to all abstract limiting procedures. In this sense, the most general context of integration theory is one where we have an algebraic object, a vector space for example, endowed with an operation which can be called "limit", and which satisfies some mild restrictions related to the everyday use of such objects. This is the context of general convergence structures, on algebraic objects, in the sense of Dolecki and Mynard (2016), and recently studied by Bilokopytov et al. (2024) and O’Brien et al. (2023), but already present in some aspects in Higgs (1988).
By using these ideas and the fact that the important element in every integration procedure is the way we set approximating nets of varying objects (such as sets and functions or/and combinations of them), the main proposals of integration procedures in this work, which are called the (abstract) net Riemann, abstract Lebesgue and Saks integrals, contained in Definitions 2.4, 5.1 and 6.1, are constructed in full generality with respect to their characteristic approximation procedures, generalizing earlier work on the Lebesgue case by Boccuto and Dimitriou (2015), Boccuto and Sambucini (1998b), Ballvé and Jiménez Guerra (1992b). Also, the net Riemann, our main focus, and Saks integrals are, as far as we know, new is this generality, the first one being based on more restrictive constructions of Szabó and Száz (1988) and Fleischer (1987).
In this context, we prove general additivity and isotonicity of the net Riemann integral in Theorem 2.3, as well as properties having to do with subset integration in Theorem 2.4, in which only convergence related properties are used (i.e, no topology or order), and integration of simple functions, all new in this generalized framework, and which are established in a way such that most, or all, known integration procedures satisfies them. We also develop, using elements of duality of convergence vector spaces as in Beattie and Butzmann (2002), a very general and complete Pettis integration procedure, and prove that it has a Riemmanian representation in terms of the net Riemann integral in Theorem 2.5, thus generalizing recent results in the stochastic integration literature as in Benth and Galimberti (2022), and the classical results of Sion (1973). The same is done for the Lebesgue and Saks constructions in Sections 5.1, 5.2 and 6.
As a byproduct of such general definitions and properties, we show on Section 2.1 that most of the Riemann type constructions of integrals in the literature, such as the classical Henstock-Kurzweil integral in division spaces of Henstock (1991) and the refinement integrals of Kolmogoroff (1930) and Sion (1973), are special cases of our constructions. On the same section, we also provide a review of these integration procedures in a very general context, and make a list of well-known and also non-classical integration procedures that are contained in our framework, thus showing the (initial, as we shall see) potential of such ideas to unify and generalize integration methods contained in the literature.
Nevertheless these general constructions, some examples presented in the classical literature of integration theory exemplified by Kolmogoroff (1930), Goguadze (1979), Deleanu (1957) and Ionescu Tulcea (1954a), are not contained in them, and necessitate integrals based on infinite Riemann sums. Using the initial attempt of Higgs (1988) as an inspiration, we develop, in Section 3, an abstract summability (infinite series) theory on vector spaces equipped with net convergence structures of the sort cited above, and show not only that the classical dichotomy of unconditional/conditional summability and properties of these series are valid in our abstract work, but also that a Riemann type-integral based on infinite sums, called the $S^{*}$-integral, can be defined in this setting with its usual properties, and which contains the others cited. Relations with the net Riemann integral are also proved.
To obtain further properties of all these abstract integration procedures, and connect them with more commonly used ones, we also have aimed to prove general continuity of these integrals in terms of different convergence results. Following the usual schemes in the literature, we first separate these theorems in two types: the ones based on order, and the other on topology. For the first case, we obtain a sweeping abstraction of the Lebesgue measure Kurzweil integral convergence theorem of Monteiro and Fernandez (2012), contained in our Theorem 4.1, showing that it is valid for a very general version of the net Riemann integral based on non-negative set functions with values in Riesz Spaces. Similar, we obtain an outside of null sets convergence result for the $S^{*}$, and a restricted version of the net Riemann, integrals that are not contained in the previous literature, and allows us to connect these integration procedures with the Lebesgue construction of Pavlakos (1983) and others presented in Section 5.1, which also presents a wide list of integration procedures of such (Lebesgue) type that can be compared to ours in the partially ordered case. In the same context of Riesz spaces, we also obtain very general monotone and dominated convergence theorems for the net Riemann integrals, getting explicit restrictions of our general integration procedure that allows us to obtain these theorems. Similarly, in the topological case, we prove a type of (Egorov) quasi-uniform convergence theorem in the context of the net Riemann integral, generalizing earlier work of Sion (1973), Trombetta (2004) and Millington (1975), and comment on further results contained in the literature.
In the same backdrop of continuity of abstract integration, we prove two further results that are of special interest: the Henstock lemma in Theorem 4.6 and the general convergence theorem in Theorem 4.14. Both of these are proved for functions and measures with values in S-metric, or lattice normed, spaces, a generalization of topological and partially ordered Riesz spaces with general convergences. These are, as far as we can tell, not present in any other work of integration theory, and manifest the fact that convergence theorems can be obtained even if we have no topology or order involved in our integration procedure, and contains seeds for further and more general convergence theorems of this non topological/order type.
Finally, we make connections between the main integral archetypes presented, and specific examples in the literature. In fact, we solve a series of open problems in the comparison of integration procedures that were left open in the literature, from which we highlight the following:
  1. (i)
    In the AMS review (MR0450503) of Sion (1973) book, Z. Lipecki cited the open problem of comparing Sion integral with the non-locally convex construction of Turpin (1975). In Section 5.2.1, we show that the more general integral of Drewnowski and Labuda (2013) is a special case of Sion integral, thus answering Lipecki’s question.
  2. (ii)
    In a similar context, Pavlakos (1983) and Popescu (1970a) both left open the question of comparing Lebesgue integration on partially ordered spaces with Riemannian constructions in the same setting. We do this comparison in Section 5.1 using the convergence theorems cited above. In particular, we prove that both integrals are a special case of the $S^{*}$ integral in the context of infinite summability, which uses in the proof Theorem 4.1, not obtained previously in the literature.
  3. (iii)
    In his book Sion (1973), M. Sion left open the question of comparing a version of his truncation/refinement integral with the $S^{*}$ integral. We show in Section 3 that in the topological and partially ordered cases, Sion’s procedure is more general, and contains the refereed infinite Riemann sums integral.
  4. (iv)
    Lastly, Masse (1979) left open the question of comparing his semigroup submeasure integral with Sion (1973) integration procedure. We solve this problem partially in Theorem 5.7, showing that in the presence of a control measure in a metric linear space, the two integrals are equivalent.
We finish the work by showing that our general definition of the Saks integration procedure contains not only the net Riemann integral, and therefore all of the integration theories of above, but also incorporates non-additive integration procedures such as the ones in Boccuto and Riečan (2009). A list of open problems follow, which shows that further work on our general integration procedure can lead to interesting new ideas not only on measure and integration, but also on topology and related areas.
We now pass to some formal comments on references and notation.
For information about the topological, algebraic and measure-theoretic concepts that will be used in the context of integration theory in this article, we cite, for the reader convenience, McShane (1953), Fremlin (1974), Vulikh (1967), Luxemburg and Zaanen (1971), Zaanen (1983), Gähler (1977), Gähler (1978), Boccuto (1995), Boccuto and Sambucini (2023), Boccuto and Candeloro (2009a) and, specially for integration, the more recent publications of Jiang et al. (2024), de Jeu and Jiang (2022a) and de Jeu and Jiang (2022b).
@@
In in this work, we will, in general, denote spaces of any type with uppercase Latin letters $X,Y,Z$, or $\mathcal{X}$, $\mathcal{Y}$, $\mathcal{Z}$ when there is duplicity. As a rule, all partial orders in the refereed spaces, when they have one, will be denoted by $\leq$, following the custom of classical and recent works on spaces equipped with partial orders, as Vulikh (1967), Luxemburg and Zaanen (1971), Popov and Randrianantoanina (2013), and Kalauch and van Gaans (2019), as we will not work with more than one partial order, in a fixed space, at a time (therefore, minimizing the possibility of confusion). Differentiation between order notations will be made only when necessary. As usual, we use the notation $\max(x,-x)=x\vee-x=|x|,x^{+}=\max(x,0),x^{-}=\max(-x,0)$ for the, in order, absolute value, positive and negative parts of an element $x$ of a Riesz space (vector lattice). It will also be supposed that every partially ordered space below is Archimedean, and that every topological space is Hausdorff, and we only cite explicitly these properties again in case we need to stress their presence - for more details on these notions, see Luxemburg and Zaanen (1971) and Kelley (2017). Finally, all algebraic operations shall be supposed to be commutative, and all vector spaces will be assume to be real (i.e, vector spaces with respect to the field $\mathbb{R}$).
@@
With regards to notation, for a general set $A$, following Bourbaki (2013), we denote by $\mathcal{F}(A)$ the collection of all its finite subsets (which is a directed set by inclusion $\subseteq$). We also have the following set-theoretic definition (see, for example, Rao (2018)):
@@
We say that a pair $(\Omega,\mathcal{H})$ consisting of an arbitrary (non-empty) set and a collection of its subsets is a Paved Space if $\emptyset\in\mathcal{H}$.
For the measure and integration results below, we shall need a variety of different convergence notions used in the literature, sometimes not equivalently defined in all works, and even not equivalent in general. In this case, comparison of our results with the literature should be done considering in detail the definitions below. We usually adopt the most general one to avoid problems of compatibility, and point out equivalences (and differences) wherever possible. As the topological notions are fairly unified (see, for example, the classics by Bourbaki (2013) and Kelley (2017)), we focus on the partially ordered spaces case, for which convergences are more numerous in the literature.
1.1 Partially Ordered Spaces and Order Convergence.
Considering the context above, we begin by the following definition given in Pavlakos (1983) and Kalauch and van Gaans (2019), which originally is due to Vulikh (1967), called $(o_{3})$-convergence:
@@
Let $X$ be a partially ordered set and II, $C$, and $D$ directed sets. We say that a net $\{x_{i}\}_{i\in I}$ in $X$ $(o_{3})$-converges to $x$ in $X$, which we denote by:
@@
if there are directed sets $C,D$ and an increasing net $\{z_{c}\}_{c\in C}$ and a decreasing net $\{y_{d}\}_{d\in D}$ in $X$ such that:
@@
Sometimes we may get a simplification by obtaining increasing and decreasing nets in the definition above with same indexing directed set, and then get the following definition of O’Brien et al. (2023) and Bilokopytov et al. (2024) (see also Kalauch and van Gaans (2019)):
@@
Let $X$ be a partially ordered set and $I$ a directed set. We say that a net $\{x_{i}\}_{i\in I}$ in $X$ order converges, or $(o)$-converges, to $x$ in $X$, which we denote by:
@@
if there is a directed set $C$ and an increasing net $\{z_{c}\}_{c\in C}$ and a decreasing net $\{y_{c}\}_{c\in C}$ in $X$ such that:
@@
Definition 1.4 (Bounded Nets and Absolute value of Functions).
Let $X$ be a Riesz Space and $\{x_{i}\}_{i\in I}$ a net in $X$. We say that this net is bounded if and only if there exists an $y\in X$ such that:
@@
Also, for a function $f:\Omega\rightarrow X$, where $\Omega$ is a non-empty set, we shall denote by $|f|:\Omega\rightarrow X^{+}$ the function given by:
$|f|(\omega)=|f(\omega)|,$
for each $\omega\in\Omega$.
The following characterization of order convergent nets will be useful later (see Monteiro and Fernandez (2012) and McGill (1975)):
@@
Let $X$ be a Riesz Space. Then, a net $\{x_{i}\}_{i\in I}$ is $(o)$-convergent to $x\in X$ if and only if it is bounded and:
@@
For Riesz Spaces, we have the following result of Pavlakos (1983) and Papangelou (1964), which is also given in O’Brien et al. (2023) and Bilokopytov et al. (2024) for Riesz spaces:
@@
Let $X$ be a Riesz Space and $(I,\alpha)$ a directed set. A net $\{x_{i}\}_{i\in I}$ is $(o)$-convergent in $X$ if and only if it is $o_{3}$-convergent, which occurs if and only if there is a decreasing net $\{y_{d}\}_{d\in D}$ in $X$, with $D$ a directed set, such that $y_{d}\downarrow 0$ and for every $d\in D$ there exists $i^{*}\in I$ such that:
@@
Most of the time we will work with objects less general than lattice ordered groups, e.g. Riesz spaces, in such a way that all of the previous convergence definitions, including the one given in the proposition above, are equivalent. Also, we can show, at least for sequences, that the following result is true (see (Kalauch and van Gaans, 2019, Section 3.7, p. 184)):
@@
Let $X$ be a Riesz Space. A sequence $\{x_{n}\}_{n\in\mathbb{N}}$ in $X$ is $(o_{3}),o$-convergent to $x$ in $X$ if and only if there is a non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ in $X$ such that $u_{n}\downarrow 0$ and:
@@
The sequential nature of the result above is essential, as "tail" type convergence results for nets are not usually not true. For more about this subject, see Section 3.7 of Kalauch and van Gaans (2019).
For various proofs in the analytical theory of partially ordered spaces, one hard to work with property of the previous convergences is that $\varepsilon$-type arguments have no analogue. This complicates some proofs related to Cauchy-type conditions and countable sequence and subsequence arguments. To remediate this fact, a different type of convergence, usually spoken as a type of "regulator" or $(D)$-convergence was created, and plays a major role in classical texts, such as Zaanen (1983) and Luxemburg and Zaanen (1971), as well as in the vast literature in integration and convergence theory in partially ordered spaces, as given by the surveys of Boccuto and Dimitriou (2015), Boccuto et al. (2012b), and the recent works of O’Brien et al. (2023) and Bilokopytov et al. (2024), and their collaborators.
In this case we introduce the following structure associated with $(D)$-convergence, as in Boccuto et al. (2012b):
@@
Let $X$ be a Riesz space. We say that a double sequence $\{a_{ij}\}_{(i,j)\in\mathbb{N}\times\mathbb{N}}$ [^1] From now on, except when strictly needed, we will represent a regulator only as $\{a_{ij}\}_{ij}$ or $\{a_{ij}\}$. in $X$ is a $(D)$-sequence, or regulator, if $a_{ij}\geq a_{ij+1}$ for every $i,n\in\mathbb{N}$ and:
@@
Let $X$ be a Riesz Space and $I$ a directed set. We say that a net $\{x_{i}\}_{i\in I}$ in $X$ is $(D)$-convergent to an element $x\in X$ if there exists a $(D)$-sequence $\{a_{ij}\}$ in $X$ such that for every $\varphi\in\mathbb{N}^{\mathbb{N}}$ there exists $i_{0}\in I$ such that:
@@
In general, we have the following relations between this convergence and the previously defined ones (see Chapter 2 of Boccuto et al. (2012b)):
@@
Let $X$ be an Riesz Space. Then, a $(o)$-convergent sequence $\{x_{n}\}_{n\in\mathbb{N}}$ in $X$ is $(D)$-convergent.
In general, the reverse statement of convergence is not true - see, for example, Luxemburg and Zaanen (1971). The result for nets is also not, in general, true - see again Chapter 2 of Boccuto et al. (2012b).
As indicated earlier, $(D)$-convergence can give, in some cases, a substitute for $\varepsilon\setminus 2^{n}$-type arguments in terms of (one version of) the Fremlin Lemma (see Riečan and Neubrunn (1997)), for which we need the next notions, beginning with the essential concept from the integration theory of Matthes (1961) and Maitland-Wright (1971).
Definition 1.7 (Weak $\sigma$-Distributivity).
Let $X$ be a Riesz Space. We say that $X$ is weakly $\sigma$-distributive if and only if:
for every $(D)$-sequence $\{a_{ij}\}$ in $X$.
As an interesting consequence (which we shall use later in the context of integration) of this property, we have (see (Boccuto et al., 2012b, Proposition 2.22, p. 31)):
Proposition 1.5.
Let $X$ be an Archimedean and weakly $\sigma$-distributive Riesz Space. Then, a $(D)$-convergent sequence $\{x_{n}\}_{n\in\mathbb{N}}$ in $X$ is $(o)$-convergent.
The last definition we shall need for the Fremlin Lemma is the next classical concept, which can be found in most books about partially ordered spaces (see, for example, Vulikh (1967) or Luxemburg and Zaanen (1971)):
Definition 1.8 (Dedekind Completeness).
For further properties relating to Dedekind completeness, see Chapter 4 of Luxemburg and Zaanen (1971).
We now have the main lemma we alluded to, and which finishes this section:
@@
A detailed proof of this important result can be found in monteirodissertação.
1.2 Lattice Normed Spaces.
For the development of our integration theory, besides the partially ordered and topological cases, we shall also need some basic elements of convergence in spaces equipped with abstract distance functions that are valued in (a collection of) partially ordered vector spaces. This idea has been considered in various forms in the literature, and we shall drawn mainly from Kusraev (2003) and Kusraev and Malyugin (1988), but generalizing the construction to cover non-locally convex and other types of spaces. A similar notion in the topological literature is the category of S-metric spaces, studied mainly by Reichel and Ruppert (1977) and Reichel (2006).
We begin with the following definition:
Definition 1.9 (Lattice Norm).
Let $X$ be a (real) vector space and $\mathcal{X}$ a Riesz space and $\mathcal{X}^{+}$ its positive cone. Then, a mapping $||\cdot||:X\rightarrow\mathcal{X}^{+}$ is a lattice norm of $X$ with values in $\mathcal{X}$ if it satisfies the following for every $x,y\in X$:
  1. (i)
    $x||=0\iff x=0$
  2. (ii)
[…]
With this, we now define the main structure associated to a lattice norm:
Definition 1.10 (Lattice Normed Space).
We say that a triple $(X,\{||\cdot||_{l\in L},\{\mathcal{X}\}_{l\in L})$, where $L$ is an index set, $||\cdot||_{l}:X\rightarrow\mathcal{X}^{+}_{l}$ a lattice norm for every $l\in L$ and $\mathcal{X}_{l}$ a Riesz space for every $l\in L$, is a lattice normed space, or only $X$ if no confusion arises.
As in Chapter 2 of Kusraev (2003), this construction wields a natural notion of convergence in $X$:
Definition 1.11 (Convergence in Lattice Normed Spaces).
We say that a net $\{x_{i}\}_{i}$ of elements of a lattice normed space $(X,\{||\cdot||_{l\in L},\{\mathcal{X}\}_{l\in L})$ is convergent to an element $x\in X$ if and only if for each $l\in L$:
$x-x_{i}||_{l}\rightarrow 0,$
where this convergence is $(o)$ or $(D)$-convergence in $\mathcal{X}_{l}$ [^2] Any other net convergence structure is admitted as well, see the next section..
We then say that the lattice norm $||\cdot||_{l}$ is continuous if:
[…]
with $\lim$ being taken in the sense above. This convergence will also be said to be complete if any Cauchy net for this convergence is convergent in XX [^3] See also the next section for more details about Cauchy concepts in general convergences..
For the fact that these two constructions generalize topological as well as Riesz space convergences, we refer the reader to Kusraev (2003) and the paper Reichel (2006).
We shall now pass to a discussion of a general structure that unifies all the presented convergence concepts so far.
1.3 Net Convergence Structures in General Spaces.
To unify the different notions of order-based and topological convergences above, we shall need a concept that abstracts the notion of convergence independently of the presence of a topology or order in the space in question. Various notions of this idea exists, with varying generality, in the works Gähler (1977), Gähler (1978), Dolecki (2024), Dolecki and Mynard (2016), Higgs (1988) and others. Usually, this notion of "convergence structure" is presented in the language of filters, which are more adapted to topology. As our focus is integration, nets are more intuitive and useful. Therefore, we shall use the recent contributions of Bilokopytov et al. (2024) and O’Brien et al. (2023), which developed this theory in the context of nets.